OpenStudy (anonymous):

y= e^x^3 find y'(1) a. 3e^2 b. 3e c. e d. 1/e e. e^3

OpenStudy (anonymous):

y= e^x^3 find y'(1) y '= (3x^2)(e^x^3 ) y '(1)=3e ans

OpenStudy (anonymous):

b) 3e